# MedQARo: A Large-Scale Benchmark for Medical Question Answering in Romanian

### v2

@@ Abstract
Question answering (QA) is an actively studied topic, being a core natural language processing (NLP) task that needs to be addressed before achieving Artificial General Intelligence (AGI). However, the lack of QA datasets in specific domains and languages hinders the development of robust AI models able to generalize across various domains and languages. To this end, we introduce MedQARo, the first large-scale medical QA benchmark in Romanian, alongside a comprehensive evaluation of state-of-the-art large language models (LLMs). We construct a high-quality and large-scale dataset comprising 102,646 QA pairs related to cancer patients. The questions regard medical case summaries of 1,011 patients, requiring either keyword extraction or reasoning to be answered correctly. MedQARo is the result of a time-consuming manual annotation process carried out by seven physicians specialized in oncology or radiotherapy, who spent a total of about 2,100 work hours to generate the QA pairs. We experiment with four LLMs from distinct families of models on MedQARo. Each model is employed in two scenarios, namely one based on zero-shot prompting and one based on supervised fine-tuning. Our results show that fine-tuned models significantly outperform their zero-shot counterparts, clearly indicating that pretrained models fail to generalize on MedQARo. Our findings demonstrate the importance of both domain-specific and language-specific fine-tuning for reliable clinical QA in Romanian. We publicly release our dataset and code at https://github.com/ana-rogoz/MedQARo.
@@
1 Introduction
The development of robust question answering (QA) systems is an important goal in natural language processing (NLP), representing one of prerequisites for building systems that reach human-level intelligence. Despite the significant advances in the area of Large Language Models (LLMs) Pal and Sankarasubbu (2024); Touvron et al. (2023); Jiang et al. (2023); Taori et al. (2023); Köpf et al. (2023), which are able to perform remarkably well across various QA tasks Brown et al. (2020); Chowdhery et al. (2023); Rajpurkar et al. (2016); Hendrycks et al. (2021); Achiam et al. (2023); Singhal et al. (2023, 2025), QA remains a challenging task within specialized domains, such as the medical field, or in low-resource languages, such as Romanian. For instance, the task of medical QA comes with unique challenges due to the complexity of clinical narratives and the necessity for deep understanding of medical terms. Moreover, there is a soaring demand for high accuracy, as mistakes can have serious consequences in daily medical practice assisted by AI agents. While large-scale datasets such as PubMedQA Jin et al. (2019), emrQA Pampari et al. (2018), and MedQA Jin et al. (2021) have driven significant progress in English medical QA, the availability of similar resources in other languages, especially underrepresented ones, remains limited. To the best of our knowledge, for the Romanian language, there are no prior efforts in the realm of large-scale QA resources for the medical domain.
@@ 1 Introduction
To address this gap, we introduce MedQARo, the first large-scale benchmark for Medical Question Answering in Romanian (MedQARo). As illustrated in Figure 1, the benchmark creation involves two phases, (i) manual data collection and annotation, and (ii) model benchmarking. The dataset comprises 102,646 high-quality QA pairs from real-world clinical records of 1,011 oncology patients (796 patients with breast cancer and 215 patients with lung cancer). The QA pairs are the results of a manual annotation process carried out by physicians specialized in oncology and radiotherapy, who collectively spent nearly 2,100 hours to generate the pairs, ensuring both linguistic quality and medical correctness. MedQARo includes 76,416 QA pairs about breast cancer patients and 26,230 about lung cancer patients, with questions grounded in medical case summaries (epicrises).
Dataset Language Domain # QA pairs
SQuAD Rajpurkar et al. (2016) English General knowledge 107,785
Natural Questions Kwiatkowski et al. (2019) English General knowledge 307,373
XQuAD Artetxe et al. (2020) Multilingual (11 languages) General knowledge 1,190
MLQA Lewis et al. (2020) Multilingual (7 languages) General knowledge 17,000
TyDi QA Clark et al. (2020) Multilingual (11 languages) Information-seeking 204,000
PubMedQA Jin et al. (2019) English Medical 212,300
emrQA Pampari et al. (2018) English Medical 455,837
MedQA Jin et al. (2021) English Medical exams 12,723
MedMCQA Pal et al. (2022) English Medical exams (India) 193,100
JuRo Crăciun et al. (2025) Romanian Legal 10,836
RoITD Nicolae and Tufiş (2021) Romanian Technical 9,500
LiRo / XQuAD-Ro Dumitrescu et al. (2021) Romanian General knowledge 1,190
RoMedQA Dima et al. (2024) Romanian Biology exams 4,127
MedQARo (ours) Romanian Medical cases 102,646
@@
To prevent data leakage across splits and support fair model development and evaluation, we provide official training, validation, and test splits at the patient level, i.e. a medical record can appear in only one of the splits. We benchmark four models on MedQARo, namely two LLMs specialized on Romanian (RoLLaMA2-7B and RoMistral-7B) Masala et al. (2024), a long-context LLM (Phi-4-mini-instruct) Abouelenin et al. (2025), and an LLM tuned on biomedical data (LLaMA3-OpenBioLLM-8B) Pal and Sankarasubbu (2024). Each model is assessed in both zero-shot and supervised fine-tuning setups, while testing various prompt formats. Our results show that all LLMs significantly benefit from fine-tuning, whereas zero-shot variants perform poorly. These findings emphasize the critical importance of both domain and language adaptation for reliable clinical QA in Romanian. Among all evaluated models, the fine-tuned Phi-4-mini-instruct achieves the best performance, significantly outperforming other approaches. However, the best-performing model achieves an F1 score of only $0.667$, suggesting that MedQARo is a challenging dataset, highlighting the need for the development of more robust models, capable of better generalizing to specialized domains and low-resource languages.
In summary, our contributions are fourfold:
@@
  • We introduce MedQARo, the first large-scale medical QA dataset in Romanian, comprising 102,646 QA instances obtained with the help of trained physicians.
@@
  • We benchmark four LLMs from distinct families of models across multiple configurations and report comprehensive performance metrics under both zero-shot and fine-tuning configurations.
@@ 2 Related Work
English and multi-lingual QA datasets. Question answering emerged as a fundamental task in NLP, with several datasets being developed to support research across various domains and languages. Popular QA datasets focusing on the English language, such as SQuAD Rajpurkar et al. (2016) and Natural Questions Kwiatkowski et al. (2019), established the foundation for extractive and generative QA systems. These datasets showed the importance of large-scale high-quality annotations for training robust QA models. Later on, given the need for cross-language understanding, more and more datasets covering several languages started to arise. Some notable multilingual efforts include XQuAD Artetxe et al. (2020), which extends SQuAD to 11 languages, and MLQA Lewis et al. (2020), which provides QA datasets in seven languages with cross-lingual evaluation capabilities. Furthermore, TyDi QA Clark et al. (2020) advanced multilingual QA by covering structurally different languages and focusing on information-seeking questions. However, these efforts did not include Romanian among the target languages.
Medical QA datasets. Medical QA represents a challenging domain that requires deep understanding of a broad range of aspects, such as clinical terminology and medical reasoning. Several English medical QA datasets have been already developed to tackle the particularities of the medical domain. For instance, MedQA Jin et al. (2021) provides over 12,000 multiple-choice questions from medical licensing exams across three countries, testing complex clinical reasoning abilities. Additionally, the emrQA dataset Pampari et al. (2018) introduces one of the largest QA medical resources with over 455,000 question-answer pairs automatically generated from electronic medical records. PubMedQA Jin et al. (2019) focuses on biomedical literature comprehension, with questions derived from research article titles and abstracts. MedMCQA Pal et al. (2022) expands medical QA to include Indian medical entrance exams, with 193,100 questions across 21 medical subjects, demonstrating the importance of diverse medical knowledge sources. Despite its data source, the QA pairs contained by MedMCQA are in English.
@@
Romanian QA datasets. For the Romanian language, a few QA datasets have been independently developed. Two such datasets focus on domain specific understanding. One is the JuRo dataset Crăciun et al. (2025), which provides a collection of Romanian juridical texts with associated questions, and the other is the RoITD dataset Nicolae and Tufiş (2021), which addresses technical domain questions in the IT sector. Additionally, the LiRo benchmark Dumitrescu et al. (2021), which provides a comprehensive evaluation platform for Romanian NLP tasks, contributed with a subset of XQuAD Artetxe et al. (2020) for Romanian, consisting of translated versions of English questions. For advanced biology, Dima et al. (2024) published a small-scale dataset comprising 4,127 single-choice questions. Existing datasets for Romanian QA are typically small (containing between 1,000 and 11,000 QA pairs). With 102,646 QA pairs, MedQARo is one order of magnitude larger than the largest dataset for Romanian QA.
@@
MedQARo vs. QA dataset landscape. In Table 1, we present a comprehensive comparison between MedQARo and existing QA datasets across different languages and domains. MedQARo fills in a gap in the landscape of medical QA resources by providing the first large-scale Romanian medical QA dataset with 102,646 question-answer pairs focused on oncology. While other Romanian QA datasets exist, none of them includes extractive and reasoning questions based on medical case summaries. This positions MedQARo as the first large-scale resource for advancing medical NLP capabilities for Romanian, an underrepresented language in the QA field.
Number of patients Number of samples
@@
Lung cancer 150 32 33 215 18,300 3,904 4,026 26230
Total 707 151 153 1,011 71,772 15,328 15,546 102,646
Table 2: A summary of the number of patients and corresponding QA pairs included in the MedQARo benchmark across two types of cancer, breast and lung. The splits are performed at the patient level, into training (70%), validation (15%), and test (15%), to prevent information leakage across splits. We report the number of patients per cancer type per split, and the corresponding number of QA samples generated for each split. The final row aggregates the total number of unique patients and QA pairs across both cancer types for each split. MedQARo comprises a total 102,646 QA pairs for 1,011 patients.
3 Dataset
3.1 Data Collection
MedQARo is obtained through the manual extraction of information from medical documents, namely medical summaries of patients registered in the Oncology Department of Colţea Clinical Hospital in Bucharest, Romania. The epicrises, originally drafted in Word format by the physicians attending each patient, were manually reviewed by a group of seven physicians involved in constructing MedQARo. In this group, two experienced physicians, who led the data collection process, established the most prevalent cancer sites in their hospital, namely breast and lung. They selected patients with either one of these diagnostics, who received treatment at the hospital between 2019 and 2024. This selection process led to a set of 796 patients with breast cancer and 215 patients with lung cancer.
The two experienced physicians further proposed a set of reference questions for each cancer site, comprising 48 questions for breast cancer and 61 questions for lung cancer. The other five physicians proposed 8 to 10 equivalent reformulations of the reference questions. Next, the physicians proceeded by extracting the answers for each unique question from each epicrisis and entering the respective answers into a database.
Data extraction and completion were carried out via a two-level process. The first level was performed by five medical oncology and radiotherapy residents, who meticulously completed the database of QA pairs following prior training based on the annotation guidelines provided by the experienced physicians. The second level was carried out by two board-certified medical oncology specialists, holders of a good standing certificate issued by the Romanian College of Physicians and with proven experience in good medical and clinical research practices. They verified the accuracy and consistency of the data entered by the residents. In total, the annotation process involved seven physicians, with a cumulative effort of approximately 2,100 hours (equivalent to about 300 hours per physician). The number of patients processed by each resident ranged from a minimum of 200 to a maximum of 220 cases, depending on individual availability and case complexity. To measure the inter-annotator agreement, a number of 350 QA pairs were annotated by all five physicians. The average Cohen’s $\kappa$ score between each pair of annotators is $0.9562$, which indicates an almost perfect agreement.
Furthermore, each experienced specialist meticulously reviewed the QA pairs associated with about 500-510 patients. This distribution of the workload was planned to ensure a relatively balanced annotation process. For data extraction, the experienced physicians developed an annotation guideline, establishing clear completion rules, including: staging of cancers according to the Tumor Node Metastasis (TNM) classification; grading of adverse events according to the Common Terminology Criteria for Adverse Events (CTCAE) system; standardization of medical abbreviations (e.g. “CHT” for chemotherapy, “RT” for radiotherapy, “NK” for “not known” values); handling incomplete information. In addition to these general rules, the guideline included a detailed list of answer options to be extracted and recorded in a standardized format: Eastern Cooperative Oncology Group (ECOG) performance status, menopausal status, molecular subtype, tumor proliferation index (Ki-67), histological grade, tumor size, family history of cancer, personal history of other cancers, body mass index (BMI), tumor markers, presence/absence of disease progression, number of metastatic sites, metastatic locations, death, etc.
Regarding research ethics, we emphasize that all data was fully anonymized prior to processing by removing any direct identifiers (names, dates of birth, personal identification numbers, etc.). The use of anonymized data for research purposes is based on the informed consent of patients, as recorded in the medical chart of Colţea Clinical Hospital and signed individually by each patient at admission. The database was approved by the Ethics Committee of Colţea Clinical Hospital where the research was conducted, in accordance with Decision No. 19091, dated October 5th, 2021. The entire process complies with the requirements of the General Data Protection Regulation (GDPR), the principles of the Declaration of Helsinki, and the standards of good medical and research practices.

### v3

@@ Abstract
Question answering (QA) is an actively studied topic, being a core natural language processing (NLP) task that needs to be addressed before achieving Artificial General Intelligence (AGI). However, the lack of QA datasets in specific domains and languages hinders the development of robust AI models able to generalize across various domains and languages. To this end, we introduce MedQARo, the first large-scale medical QA benchmark in Romanian, alongside a comprehensive evaluation of state-of-the-art (SOTA) large language models (LLMs). We construct a high-quality and large-scale dataset comprising 105,880 QA pairs related to cancer patients from two medical centers. The questions regard medical case summaries of 1,242 patients, requiring either keyword extraction or reasoning to be answered correctly. MedQARo is the result of a time-consuming manual annotation process carried out by seven physicians specialized in oncology or radiotherapy, who spent a total of about 3,000 work hours to generate the QA pairs. Our benchmark contains both in-domain and cross-domain (cross-center and cross-cancer) test collections, enabling a precise assessment of generalization capabilities. We experiment with four open-source LLMs from distinct families of models on MedQARo. Each model is employed in two scenarios, namely one based on zero-shot prompting and one based on supervised fine-tuning. We also evaluate two state-of-the-art LLMs exposed only through APIs, namely GPT-5.2 and Gemini 3 Flash. Our results show that fine-tuned models significantly outperform zero-shot models, clearly indicating that pretrained models fail to generalize on MedQARo. Our findings demonstrate the importance of both domain-specific and language-specific fine-tuning for reliable clinical QA in Romanian. We publicly release our dataset and code at https://github.com/ana-rogoz/MedQARo.
keywords:
Medical question answering, medical resource, Romanian language resource, medical LLMs.
1 Introduction
The development of robust question answering (QA) systems is an important goal in natural language processing (NLP), representing one of the prerequisites for building systems that reach human-level intelligence. Despite the significant advances in the area of Large Language Models (LLMs) [OpenBioLLMs, touvron2023llama, jiang2023mistral7b, alpaca, kopf2024openassistant], which are able to perform remarkably well across various QA tasks [brown2020language, chowdhery2023palm, rajpurkar-etal-2016-squad, hendrycksmeasuring, openai2023gpt4, singhal2023large, singhal2023towards], QA remains a challenging task within specialized domains, such as the medical field, or in low-resource languages, such as Romanian. For instance, the task of medical QA comes with unique challenges due to the complexity of clinical narratives and the necessity for deep understanding of medical terms. Moreover, there is a soaring demand for high accuracy, as mistakes can have serious consequences in daily medical practice assisted by AI agents. While large-scale datasets such as PubMedQA [jin-etal-2019-pubmedqa], emrQA [pampari-etal-2018-emrqa], and MedQA [jin2020diseasedoespatienthave] have driven significant progress in English medical QA, the availability of similar resources in other languages, especially underrepresented ones, remains limited. This hinders the development of LLMs that are robust to language and domain shift. To the best of our knowledge, for the Romanian language, there are no prior efforts in the realm of large-scale QA resources for the medical domain.
@@ 1 Introduction
To address this gap, we introduce MedQARo, the first large-scale benchmark for Medical Question Answering in Romanian. As illustrated in Figure 1, the benchmark creation involves two phases, (i) manual data collection and annotation, and (ii) model benchmarking. The dataset comprises 105,880 high-quality QA pairs from real-world clinical records of 1,242 oncology patients (796 patients with breast cancer, 215 patients with lung cancer, and 231 patients with other cancers). The QA pairs are the results of a manual annotation process carried out by physicians specialized in oncology and radiotherapy, who collectively spent nearly 3,000 hours to generate the pairs, ensuring both linguistic quality and medical correctness. MedQARo includes 76,416 QA pairs about breast cancer patients, 26,230 about lung cancer patients and 3,234 about patients with other cancer types, with questions grounded in medical case summaries (epicrises).
@@
To prevent data leakage across splits and support fair model development and evaluation, we provide official training, validation, and test splits at the patient level, i.e. a medical record can appear in only one of the splits. Moreover, we provide both in-domain and cross-domain test collections, where the cross-domain test set comprises patients from a distinct medical center, who suffer from five cancer types (head and neck, melanoma, renal cell carcinoma, bladder cancer, hepatocellular carcinoma) that are not available during training. Hence, MedQARo enables a precise assessment of generalization capabilities across different medical centers and cancer types. Oncologists, especially residents, could greatly benefit from using LLMs to extract and reason about the information included in epicrises, when they need to recommend diagnostic tools or treatment lines for their patients. MedQARo provides the necessary means to develop such LLMs, that could be deployed in daily medical practice. We benchmark four models on MedQARo, namely two LLMs specialized on Romanian (RoLLaMA2-7B and RoMistral-7B) [masala2024vorbecstiromanecsterecipetrain], a long-context LLM (Phi-4-mini-instruct) [phi4mini], and an LLM tuned on biomedical data (LLaMA3-OpenBioLLM-8B) [OpenBioLLMs]. Each model is assessed in both zero-shot and supervised fine-tuning setups, while testing various prompt formats. In addition, we assess two state-of-the-art LLMs in the zero-shot setting, which are only accessible through paid APIs, namely GPT-5.2 and Gemini 3 Flash. Our results show that all LLMs significantly benefit from fine-tuning, whereas zero-shot variants perform poorly. These findings emphasize the critical importance of both domain and language adaptation for reliable clinical QA in Romanian. Among all evaluated models, the fine-tuned Phi-4-mini-instruct achieves the best performance, significantly outperforming other approaches. However, the best-performing model achieves an F1 score of only $0.667$ on the in-domain test set, suggesting that MedQARo is a challenging dataset, highlighting the need for the development of more robust models, capable of better generalizing to specialized domains and low-resource languages.
In summary, our contributions are fivefold:
@@
  • We introduce MedQARo, the first large-scale medical QA dataset in Romanian, comprising 105,880 QA instances obtained with the help of trained physicians.
@@
  • We benchmark four LLMs from distinct families of models across multiple configurations and report comprehensive performance metrics under zero-shot and fine-tuning configurations, in both in-domain and cross-domain settings.
@@
  • We show that state-of-the-art API-based LLMs, such as GPT-5.2 and Gemini 3 Flash, perform far worse than smaller fine-tuned LLMs, further confirming the importance of fine-tuning.
@@ 2 Related Work
English and multi-lingual QA datasets. Question answering emerged as a fundamental task in NLP, with several datasets being developed to support research across various domains and languages. Popular QA datasets focusing on the English language, such as SQuAD [rajpurkar-etal-2016-squad] and Natural Questions [kwiatkowski-etal-2019-natural], established the foundation for extractive and generative QA systems. These datasets showed the importance of large-scale high-quality annotations for training robust QA models. Later on, given the need for cross-language understanding, more and more datasets covering several languages started to arise. Some notable multilingual efforts include XQuAD [artetxe-etal-2020-cross], which extends SQuAD to 11 languages, and MLQA [lewis-etal-2020-mlqa], which provides QA datasets in seven languages with cross-lingual evaluation capabilities. Furthermore, TyDi QA [clark-etal-2020-tydi] advanced multilingual QA by covering structurally different languages and focusing on information-seeking questions. However, these efforts did not include Romanian among the target languages.
Dataset Language Domain # QA pairs
SQuAD [rajpurkar-etal-2016-squad] English General knowledge 107,785
Natural Questions [kwiatkowski-etal-2019-natural] English General knowledge 307,373
XQuAD [artetxe-etal-2020-cross] Multilingual (11 languages) General knowledge 1,190
MLQA [lewis-etal-2020-mlqa] Multilingual (7 languages) General knowledge 17,000
TyDi QA [clark-etal-2020-tydi] Multilingual (11 languages) Information-seeking 204,000
PubMedQA [jin-etal-2019-pubmedqa] English Medical 212,300
emrQA [pampari-etal-2018-emrqa] English Clinical cases 455,837
MedQA [jin2020diseasedoespatienthave] English Medical exams 12,723
MedMCQA [pmlr-v174-pal22a] English Medical exams (India) 193,100
JuRo [craciun-etal-2025-graf] Romanian Legal 10,836
RoITD [nicolae2021roitd] Romanian Technical 9,500
LiRo / XQuAD-Ro [liro2021] Romanian General knowledge 1,190
RoMedQA [dima-etal-2024-roqllama] Romanian Biology exams 4,127
MedQARo (ours) Romanian Clinical cases 105,880
Medical QA datasets. Medical QA represents a challenging domain that requires deep understanding of a broad range of aspects, such as clinical terminology and medical reasoning. Several English medical QA datasets have been already developed to tackle the particularities of the medical domain. For instance, MedQA [jin2020diseasedoespatienthave] provides over 12,000 multiple-choice questions from medical licensing exams across three countries, testing complex clinical reasoning abilities. Additionally, the emrQA dataset [pampari-etal-2018-emrqa] introduces one of the largest QA medical resources with over 455,000 question-answer pairs automatically generated from electronic medical records. PubMedQA [jin-etal-2019-pubmedqa] focuses on biomedical literature comprehension, with questions derived from research article titles and abstracts. MedMCQA [pmlr-v174-pal22a] expands medical QA to include Indian medical entrance exams, with 193,100 questions across 21 medical subjects, demonstrating the importance of diverse medical knowledge sources. Despite its data source (Indian medical exams), the QA pairs contained by MedMCQA are in English.
@@
Romanian QA datasets. For the Romanian language, a few QA datasets have been independently developed. Two such datasets focus on domain specific understanding. One is the JuRo dataset [craciun-etal-2025-graf], which provides a collection of Romanian juridical texts with associated questions, and the other is the RoITD dataset [nicolae2021roitd], which addresses technical domain questions in the IT sector. Additionally, the LiRo benchmark [liro2021], which provides a comprehensive evaluation platform for Romanian NLP tasks, contributed with a subset of XQuAD [artetxe-etal-2020-cross] for Romanian, consisting of translated versions of English questions. For advanced biology, dima-etal-2024-roqllama published a small-scale dataset comprising 4,127 single-choice questions. Existing datasets for Romanian QA are typically small (containing between 1,000 and 11,000 QA pairs). With 105,880 QA pairs, MedQARo is one order of magnitude larger than the largest dataset for Romanian QA.
@@
MedQARo vs. QA dataset landscape. In Table 1, we present a comprehensive comparison between MedQARo and existing QA datasets across different languages and domains. MedQARo fills in a gap in the landscape of medical QA resources by providing the first large-scale Romanian medical QA dataset with 105,880 question-answer pairs focused on oncology. While other Romanian QA datasets exist, none of them includes extractive and reasoning questions based on medical case summaries. This positions MedQARo as the first large-scale resource for advancing medical NLP capabilities for Romanian, an underrepresented language in the QA field.
3 Dataset
3.1 Data Collection
MedQARo is obtained through the manual extraction of information from medical documents, namely medical summaries of patients that are registered either in the Oncology Department of Colţea Clinical Hospital, or in the “Dan Furtună” Medical Center, both located in Bucharest, Romania. The epicrises, originally drafted in Word format by the physicians attending each patient, were manually reviewed by a group of seven physicians involved in constructing MedQARo. In this group, two experienced physicians from Colţea Clinical Hospital, who led the data collection process, established the most prevalent cancer sites in their hospital, namely breast and lung. They selected patients with either one of these diagnoses, who received treatment at the hospital between 2019 and 2024. This selection process led to a set of 796 patients with breast cancer and 215 patients with lung cancer. To create a diverse multi-center and multi-cancer benchmark, 231 patients with five cancer types (head and neck, melanoma, renal cell carcinoma, bladder cancer, hepatocellular carcinoma) were included in the study. These patients are enrolled in cancer follow-up care programs at “Dan Furtună” Medical Center.
The two experienced physicians further proposed a set of reference questions for each cancer site, comprising 48 questions for breast cancer, 61 questions for lung cancer, and 7 general questions that cover all other types. The other five physicians proposed 8 to 10 equivalent reformulations of the reference questions. Next, the physicians proceeded by extracting the answers for each unique question from each epicrisis and entering the respective answers into a database. To optimize annotation time, the questions for each epicrisis are grouped and annotated in the same batch. This means that a physician could read an epicrisis once, before answering multiple questions about the respective patient.
Data extraction and completion were carried out via a two-level process. The first level was performed by five medical oncology and radiotherapy residents, who meticulously completed the database of QA pairs following prior training based on the annotation guidelines provided by the experienced physicians. The second level was carried out by two board-certified medical oncology specialists, holders of a good standing certificate issued by the Romanian College of Physicians and with proven experience in good medical and clinical research practices. They verified the accuracy and consistency of the data entered by the residents. In total, the annotation process involved seven physicians, with a cumulative effort of approximately 2,170 hours (equivalent to about 310 hours per physician) just to type in the answers. Neither the time required to formulate questions nor the time required to read an epicrisis before answering all questions for a certain patient is taken into account in the time measurement. While the annotation interface did not allow us to measure the extra time required to read epicrises, we estimate that this step required between 600 and 800 additional hours. The time required to formulate all question variants is estimated at 50-70 hours. Adding all numbers, we estimate that the entire annotation process took nearly 3,000 hours. The number of patients processed by each resident ranged from a minimum of 200 to a maximum of 280 cases, depending on individual availability and case complexity. To measure the inter-annotator agreement, a number of 350 randomly chosen QA pairs were annotated by all five physicians. The average Cohen’s $\kappa$ score between each pair of annotators is $0.9562$, which indicates an almost perfect agreement.
Furthermore, each experienced specialist meticulously reviewed the QA pairs associated with about 600-630 patients. This distribution of the workload was planned to ensure a relatively balanced annotation process. For data extraction, the experienced physicians developed an annotation guideline, establishing clear completion rules, including: staging of cancers according to the Tumor Node Metastasis (TNM) classification; grading of adverse events according to the Common Terminology Criteria for Adverse Events (CTCAE) system; standardization of medical abbreviations (e.g. “CHT” for chemotherapy, “RT” for radiotherapy, “NK” for “not known” values); handling incomplete information. In addition to these general rules, the guideline included a detailed list of answer options to be extracted and recorded in a standardized format: Eastern Cooperative Oncology Group (ECOG) performance status, menopausal status, molecular subtype, tumor proliferation index (Ki-67), histological grade, tumor size, family history of cancer, personal history of other cancers, body mass index (BMI), tumor markers, presence/absence of disease progression, number of metastatic sites, metastatic locations, death, etc. The senior physicians also acted as referees to solve annotation conflicts, choosing the final answer between two options based on their own high-skilled assessment.
Regarding research ethics, we emphasize that all data was fully anonymized prior to processing by removing any direct identifiers (names, dates of birth, personal identification numbers, etc.). The use of anonymized data for research purposes is based on the informed consent of patients, as recorded in the medical charts of Colţea Clinical Hospital and “Dan Furtună” Medical Center, and signed individually by each patient at admission. The database was approved by the Ethics Committee of Colţea Clinical Hospital where the research was conducted, in accordance with Decision No. 19091, dated October 5th, 2021. The database was also approved by the Ethics Committee of “Dan Furtună” Medical Center, as per Decision No. 1323, dated 11th December, 2025. The entire process complies with the requirements of the General Data Protection Regulation (GDPR), the principles of the Declaration of Helsinki, and the standards of good medical and research practices.
Table 2: A summary of the number of patients and corresponding QA pairs included in the MedQARo benchmark across cancer types. For breast and lung cancer, the splits are performed at the patient level, into training (70%), validation (15%), and in-domain test (15%), to prevent information leakage across splits. A cross-domain test set comprises a separate set of 231 patients with five cancer types (excluding breast and lung), from a different medical center. We report the number of patients per cancer type per split, and the corresponding number of QA samples generated for each split. The final row aggregates the total number of unique patients and QA pairs across different cancer types for each split. MedQARo comprises a total 105,880 QA pairs for 1,242 patients.
Patient group Number of patients Number of samples
@@ 3.1 Data Collection
Lung cancer 150 32 33 215 18,300 3,904 4,026 26,230
Other cancers - - 231 231 - - 3,234 3,234
Total 707 151 384 1,242 71,772 15,328 18,780 105,880